### v2

# Training Language Models to Explain Their Own Computations

Abstract
Can language models (LMs) learn to faithfully describe their internal computations? Are they better able to describe themselves than other models? We study the extent to which LMs’ privileged access to their own internals can be leveraged to produce new techniques for explaining their behavior. Using existing interpretability techniques as a source of ground truth, we fine-tune LMs to generate natural language descriptions of (1) the information encoded by LM features, (2) the causal structure of LMs’ internal activations, and (3) the influence of specific input tokens on LM outputs. When trained with only tens of thousands of example explanations, explainer models exhibit non-trivial generalization to new queries. This generalization appears partly attributable to explainer models’ privileged access to their own internals: using a model to explain its own computations generally works better than using a different model to explain its computations (even if the other model is significantly more capable). Our results suggest not only that LMs can learn to reliably explain their internal computations, but that such explanations offer a scalable complement to existing interpretability methods. [^1] Code and data at https://github.com/TransluceAI/introspective-interp.
Figure 1: Overview of our methods for training models to describe their own internal procedures. As a first step, we extract answers to three types of question about the target model using different interpretability methods: (A) descriptions of features of the target model’s hidden representations via an auto-interp procedure, (B) explanations about what parts of the model affect the output via activation patching, (C) explanations about what parts of the input affect the output via input ablation. As a second step, we fine-tune three explainers to produce each of the three types of explanations. (D) After fine-tuning various explainer models on various target models, we find evidence of privileged access for all three explanation types, and use this to obtain data-efficient explainers.
1 Introduction
When language models (LMs) are prompted to explain decisions—to describe what features they relied on, or what intermediate computations they performed when generating an answer—their outputs are often plausible (korbak2025chainthoughtmonitorabilitynew). But even when plausible, these explanations may not have any relationship to the internal computation that gave rise to LMs’ decisions in the first place. Indeed, work using attribution, feature visualization, and mechanistic interpretability techniques (Simonyan2013DeepIC; 10.5555/3495724.3497168; templeton2024scaling) has shown that LMs’ verbalized explanations sometimes systematically fail to describe the factors that determine their decisions (turpin2023language; chen2025reasoningmodelsdontsay; barez2025cot).
Can we fix this problem, and train LMs so that their verbalized explanations faithfully describe their internal computations? Several recent papers (binder2025looking; song2025privilegedselfaccessmattersintrospection; plunkett2025selfinterpretabilityllmscomplexinternal) have studied the related question of whether models can describe features of their own output distributions, arguing that models might be able to leverage privileged access to their own internals to do so effectively. In this paper, we are interested in an even deeper form of privileged access: whether models can learn to describe not only their outputs, but their internal representations and mechanisms, and whether privileged self-access enables them to do so better than learned explanation-generating methods derived from other models. We state our main hypothesis as follows:
The Privileged Access Hypothesis Models trained to explain their own internal computations can do so more accurately than other models trained to explain them.
To explore this hypothesis, we fine-tune various explainer models to predict three aspects of target models’ internal procedures: [^2] This paper treats the target model as fixed and only fine-tunes the explainer model. While this setup does not constitute self-explanation in the strictest sense (since the explainer’s output distribution changes), we believe it is more useful to explain pre-trained models than enforce this strict notion of self-consistency. We discuss the implications of this simplification in Section 7. For the rest of this paper, we use the term “self-explaining” in this looser sense.
  1. 1.
    Descriptions of internal features, specifically the kinds of inputs that activate them (feature descriptions; Figure˜1A).
  2. 2.
    Outcomes of interventions to internal activations (activation patching; Figure˜1B).
  3. 3.
    Descriptions of decision rules, specifically the important tokens for a decision (input ablation; Figure˜1C). We focus on a setting where recent work found that reasoning models failed to produce faithful chains-of-thought (chen2025reasoningmodelsdontsay), in order to explore whether training can improve faithfulness. This setting also bears similarities to binder2025looking due to measuring self-access to the model’s output distributions.
Our main experiments focus on feature descriptions, where we find evidence of privileged access—our target model is better explained by itself than by other explainer models, even when the other model is generally more capable (e.g. larger or instruction-tuned).
We dive deeper into this phenomenon and probe the generality of this hypothesis, finding that:
  1. 1.
    LMs can be fine-tuned to self-explain: As in past work (li2025do), we find that explanation capabilities are not always natively present in models, and must be acquired via fine-tuning. A small amount of fine-tuning enables LMs to faithfully describe their internal features, significantly outperforming zero-shot methods (kharlapenko2024self; chen2024selfie) and nearest-neighbor SAE features (huben2024sparse). Furthermore, a model’s ability to explain (its own or others’) features is correlated with the degree of similarity between the explainer and target models (§3.4).
  2. 2.
    Self-explaining is data-efficient: Self-explaining is approximately a hundred times more sample-efficient than a nearest neighbors baseline, achieving comparable results with only 0.8% of the training data (§4). This makes self-explaining a scalable approach to generating feature descriptions.
  3. 3.
    Privileged access extends to other tasks: we train models to explain the effects of activation patching (§5.1) and input ablation (§5.2). For the first task, we see the same pattern of privileged access as in feature descriptions. For the second, we see a slightly different pattern, as some models appear to be universally stronger explainers (e.g. Qwen outperforms Llama across the board); however, we still observe evidence of privileged access here (e.g. Qwen is better at explaining itself than explaining Llama, and vice versa).
Taken together, these results suggest that even when language models cannot faithfully self-explain as a result of ordinary training, they can learn to do so through an objective that enforces consistency between their external explanations and their internal procedures. This reframes interpretation as not only an external analysis problem, but as a capability that can be trained into LMs themeselves; by leveraging privileged access to internal computations, “introspective interpretability” techniques offer an avenue towards scalable understanding of model behavior.
2 Methods
In this section we describe a general framework for training models to verbalize different aspects of their internal states, and then instantiate it with three specific explanation types.
2.1 Definitions and Task
Abstractly, a neural model $\mathcal{M}$ can be represented as a pipeline $x\to h\to y$ that maps inputs $x$ to hidden representations $h$ to outputs $y$. An explanation $E$ of the neural network may enable us to answer questions $q$ about what information in $x$ is represented in $h$, how that information influences $y$, and what pieces of the model are responsible for each step.
For example, we may wish to answer a question $q=\textit{What inputs activate direction $v$ at layer $l$?}$ If $v$ activates highly on tokens Tokyo, NYC, and Paris, then an appropriate explanation $E$ might be city names. We are interested in whether LMs can learn to generate them directly.
In this paper, we will train explainer models $\mathcal{E}$ to produce explanations $E$ for various target models $\mathcal{M}$ and types of questions $q$. Abstractly, let $T:(\mathcal{M},q)\mapsto E$ be a (model-external) explanation procedure that generates an explanation $E$ answering a question $q$ about model $\mathcal{M}$ . We use $T$ as supervision to train an explainer model, by minimizing the cross-entropy loss on the explanation:
$\mathcal{L}_{\mathcal{E}}=-\mathbb{E}_{q}\left[\log p_{\mathcal{E}}(E=T(\mathcal{M},q)\mid q)\right].$ (1)
Our experiments focus on explaining Transformer language models $\mathcal{M}$ (grattafiori2024llama3herdmodels; yang2025qwen3technicalreport), which consist of layers $\ell$ each containing attention and MLP blocks, and residual stream representations $h_{\ell}$ in between each layer. We also use Transformer language models as explainers $\mathcal{E}$. Throughout this paper, we use notation $x=(x_{1},\cdots x_{n})$ to refer to an input sequence with token positions $t\in[n]$. We use $h_{(\ell,t)}(x)$ to denote the residual stream activation of $\mathcal{M}$ at layer $\ell$ and token position $t$ for input $x$.
We will train explainer LMs to answer three types of questions $q$, described below.
2.2 Feature Descriptions: What does a model-internal feature represent?
Figure 2: Training an explainer to predict feature descriptions. (A) We run a target model $\mathcal{M}$ across many input contexts. (B) We then observe vectors $v$ at layers $\ell$ and select an explanation $E$ that best describes the contexts in which $v$ is active, which we assess via a LM simulator. (C) Finally, we train an explainer model to respond with $E$ when given the question $q$ about what $v$ means at layer $\ell$, and find that it performs well on both held-out and OOD vectors $v$.
Question $q$.
For feature description, the question $q$ is what types of inputs activate direction $v$ in the residual stream at layer $l$?
Interpretability procedure $T$.
We make use of existing automated interpretability methods (hernandez2022natural; bills2023language; choi2024automatic), which generate natural-language descriptions $E$ by summarizing input tokens on which a given $v$ is highly active.
Formally, let $a_{v}(x,\ell,t)=\langle h_{\ell,t}(x),v\rangle$ denote the activation of $v$ in the residual stream representation of the token $x_{t}$. To obtain feature descriptions, we begin by prompting (or training) a “simulator” language model (LM simulator in Figure˜2) to map from candidate descriptions $E$ (e.g. Encodes city names in Figure˜2) and input sequences $x$ (e.g. we’ll go visit NYC) to predicted activations $\hat{a}(x,t,E)$ (e.g. outputting 0.1 for visit and 0.9 for NYC). See Figure˜2B. We train the simulator following choi2024automatic on FineWeb exemplars (penedo2024the) with Neuronpedia SAE labels.
Given the simulator, for each direction $v$, we search for the explanation $E$ that maximizes correlation between simulated and true activations over all input data:
$T_{\text{feat}}(\mathcal{M},v,\ell)=\operatorname*{arg\,max}_{E}\mathbb{E}_{x}\left[\texttt{corr}_{t}\left(a_{v}(x,\ell,t),\,\hat{a}(x,t,E)\right)\right]~.$ (2)
Intuitively, this procedure will assign each feature $v$ an explanation $E$ that describes the tokens on which $v$ is most active. The prompts we use to encode $q$ and $E$ are given in Appendix˜D.
Training the explainer.
$\mathcal{E}$ receives a question $q$ that contains a feature $v$ and a layer $\ell$, and is trained to generate the explanation $E$ of when $v$ is highly active at layer $\ell$. We take $v$ from the target model $\mathcal{M}$ at its layer-$\ell$ residual stream, and pass it as a continuous token to the explainer, inserted at the embedding layer. This enables the explainer to accept arbitrary vectors vv. [^3] In early experiments, we also investigated explaining features $v$ from other components of the target model, but found that explainers struggled to reproduce these explanations. We hypothesize that we need to patch features into the same type of location (i.e. the embedding layer is part of the residual stream). See Figure˜2C.
For explainer models whose hidden dimension does not match the target model, we introduce a linear projection $\Pi_{\ell}\in\mathbb{R}^{d_{\mathcal{E}}\times d_{\mathcal{M}}}$ from the target model hidden size $d_{\mathcal{M}}$ to the explainer model hidden size $d_{\mathcal{E}}$, which is trained jointly with the rest of LM parameters. We learn a separate projection per layer $\ell$ of the target model. We train explainers and projections to optimize:
$\boxed{\mathcal{L}_{\text{feat}}=-\mathbb{E}_{v,\ell}\left[\log p_{\mathcal{E}}(E=T_{\text{feat}}(\mathcal{M},v,\ell)\mid\Pi_{\ell}\,v,\ell)\right]}$ (3)
Choice of directions $v$.
Because activations $h$ tend to be highly polysemantic (elhage2022superposition), interpretation is often done on decomposed representations, which are commonly either individual neurons (so $v$ is a one-hot vector; bau2020units) or Sparse Auto-Encoder (SAEs) features obtained from sparse dictionary learning approaches (with $v$ being a member of the learned dictionary; huben2024sparse; bricken2023monosemanticity). Our experiments train on SAE features, but test generalization to full activations and activation differences. See details in Section˜3.2.
2.3 Activation Patching: What internal components affect the prediction?
Figure 3: Training an explainer to predict activation patching outcomes. (A) We run a target model $\mathcal{M}$ on an input $x$ and obtain a prediction $\mathcal{M}(x)$. (B) We perform activation patching by running $\mathcal{M}$ on a counterfactual input $x^{\prime}$—in this example, patching in vector $v$ from token $x_{1}$ and layer $\ell_{1}$ of the counterfactual run into the original run—and then construct an explanation $E$ about how the resulting prediction changes. (C) Finally, we train an explainer model to answer $E$ when given a question $q$ about the patching procedures.
Question $q$.
For activation patching, we ask which hidden components are causally important for a model’s output (meng2022locating; zhang2024towards). Specifically, the question $q$ is on input $x$, (how) does changing the activation at layer $\ell$ and token $x_{t}$ affect the output? An example is shown in Figure˜3.
Interpretability Procedure $T$.
We can answer questions of this form through activation patching and circuit tracing methods. Given an input $x$ (Paris is the capital of in Figure˜3), an activation patching experiment typically constructs an alternative input $x^{\prime}$ (Rome is the capital of). During the forward pass, we replace the activation $h_{\ell,t}(x)$ with the corresponding activation $h_{\ell,t}(x^{\prime})$:
$\mathcal{M}\left(x;\,h_{\ell,t}(x)\leftarrow h_{\ell,t}(x^{\prime})\right)~,$ (4)
and we measure the change in the model’s output:
$T_{\text{patch}}(\mathcal{M},q)=T_{\text{patch}}(\mathcal{M},x,\ell,t)=d\left(\mathcal{M}(x),\;\mathcal{M}\left(x;\,h_{\ell,t}(x)\leftarrow h_{\ell,t}(x^{\prime})\right)\right),$ (5)
where $d(\cdot,\cdot)$ is some measure of the difference between the two predictions. This process is illustrated in Figure˜3B.
Training the explainer.
Rather than directly predicting the numerical score in Equation˜5, we train explainers $\mathcal{E}$ to predict (1) whether the model output changes under patching, and (2) the content of the model output after patching. Similarly to Section˜2.2, we insert $v$ as a continuous token in the embedding layer of the explainer model. See Figure˜3C.
Additionally, rather than patching activations one layer at a time, we modify representations at several layers simultaneously to ensure sufficient samples where the model’s output changed under intervention. We divide the target model’s layers into four blocks, and when performing activation patching, compute an aggregate representation $v=\texttt{avg}_{\ell_{k}}\left(h_{\ell_{k},t}(x^{\prime})\right)$ from the counterfactual input, then insert $v$ at all corresponding layers of the original input. [^4] We perform aggregation so that we only have to specify a single continuous token to the explainer. Finally, we optimize:
$\boxed{\mathcal{L}_{\text{patch}}=-\mathbb{E}_{x,x^{\prime},t,\ell_{1\cdots i}}\left[\log p_{\mathcal{E}}\left(E=T_{\text{patch}}(\mathcal{M},x,t,\ell_{1\cdots i},v)\mid x,t,\ell_{1\cdots i},v\right)\right]}$ (6)
The exact prompt and output templates are provided in Appendix˜D.
2.4 Input Ablation: What parts of the input matter?
Figure 4: Training an explainer to predict input ablation outcomes. We run a target model $\mathcal{M}$ on an input $x$ (A), then identify whether a part of the input $\tilde{x}\subset x$ was important to the output by rerunning $\mathcal{M}$ on $\x~x\backslash\tilde${x} (B). Finally, we train an explainer model to answer question $q$ about how the output changes when $\tilde{x}$ is removed from $x$ (C).
Question $q$.
For input ablation, we ask which input tokens most influence the model’s prediction (chang-etal-2025-jopa; vafa-etal-2021-rationales). Specifically, the question $q$ is Are tokens $\tilde{x}$ in $x$ important to the model’s prediction? For instance, in Figure˜4, the question $q$ asks whether the hint is important to the answer.
Interpretability Procedure $T$.
As with activation patching, we can remove (subsets of) tokens in $x$ and identify which ones cause predictions $y$ to change. Let $\tilde{x}\subseteq x$ be a subset of tokens to remove. Input ablation measures how much the model output changes when we take an input $x$ (e.g. The first US president… in Figure˜4) and remove a subset of tokens $\tilde{x}$ (e.g. Hint: B):
[…]
where $d$ measures difference in the model’s outputs. This process is illustrated in Figure˜4B.
Training the explainer.
$\mathcal{E}$ receives a question $q$ that contains a multiple-choice reasoning question $c$ with a hint $\tilde{x}$, and is asked to predict (1) whether removing $\tilde{x}$ would change the model’s answer (the has-changed value), and (2) the new output without the hint (the content of the change). See Figure˜4C. Formally, the objective we optimize is:
$\boxed{\mathcal{L}_{\text{input}}=-\mathbb{E}_{x,\tilde{x}}\left[\log p_{\mathcal{E}}(E=T_{\text{input}}(\mathcal{M},x,\tilde{x})\mid x,\tilde{x})\right]}$ (8)
Full prompt and output templates are provided in Appendix˜D.
3 Privileged Access Improves Explanations
In this section, we (1) show that models can be fine-tuned to generate feature descriptions and generalize to new features and feature bases, and (2) provide evidence that this capability relies on privileged access. Our experiments focus on feature description (§2.2), where we train explainer models to verbalize descriptions of the target model’s features. We study additional tasks (patching, input ablation) in §5.
3.1 Models and Training Data
We study Llama-3.1-8B (grattafiori2024llama3herdmodels) as the target model, and test five different explainer models (Llama-3.1-8B, Llama-3-8B, Llama-3.1-8B-Instruct, Llama-3.1-70B, Qwen3-8B; yang2025qwen3technicalreport). [^5] Details can be found in Section A.1. We select the explainer models that differ from the target model in attributes such as architecture (Qwen3-8B), size (Llama-3.1-70B), and training regimen (Llama-3.1-8B-Instruct), to investigate the effect of each of these factors. We fine-tune each model to be an effective explainer as described in Equation˜3. For Llama-3.1-8B, we report results with both LoRA (hu2022lora) and full-parameter fine-tuning, while for Llama-3.1-70B we only report LoRA due to compute constraints, and for other models we only report full-parameter fine-tuning.
For models that require projection layers $\Pi_{\ell}$ due to hidden-size mismatch (Qwen3-8B, Llama-3.1-70B), we randomly initialize those layers and train them with the rest of the parameters. [^6] For LoRA training, these layers are included in the set of parameters trained via low-rank updates. We additionally train a version of Llama-3.1-70B where the projection layer is pre-trained to minimize $\|h^{\mathcal{E}}_{\ell,t}-\Pi_{\ell}\cdot h^{\mathcal{M}}_{\ell,t}\|_{2}$ on FineWeb samples (penedo2024the), or the distance between the the layer-$\ell$, token-$t$ explainer activation $h^{\mathcal{E}}_{\ell,t}$ , and the projected target activation $\Pi_{\ell}\cdot h^{\mathcal{M}}_{\ell,t}$ from the same layer, token location.
To generate training data, we use Llama-Scope (he2024llamascope) to obtain residual stream SAE features for Llama-3.1-8B target models, which we use for the direction $v$. We use the Llama-3.1-8B-LXR-32x features, which maps the residual stream after each layer to 131K features (32 times larger than the residual hidden dimension). [^7] We run additional experiments with Gemma 2 (gemmateam2024gemma2improvingopen), whose data generation details can be found in Section B.1. Ground-truth explanations $E$ for each feature were obtained from Neuronpedia (neuronpedia).
3.2 Evaluation Set and Metrics
We train on a subset of SAE features $S_{train}$, then evaluate generalization both in-distribution and out-of-distribution, on three types of features:
Held-out SAE features (SAE).
We hold out a subset of 1550 SAE features (50 per layer) for testing.
Full Activations (ACT).
We test out-of-distribution generalization to full residual stream activations $h_{(\ell,t)}(x)$ on inputs $x$ sampled from FineWeb (penedo2024the).
[…]
Differences between counterfactual pairs of inputs allow us to isolate the effect of a single pointwise change. We create counterfactual pairs $(x,x^{\prime})$ from the Counterfact dataset (meng2022locating) using the procedure described in Section˜2.3 and extract activation differences $v=h_{\ell,t}(x)-h_{\ell,t}(x^{\prime})$.
On each of these evaluation sets, we consider two types of metrics:
LM judge.
For SAE features, gold descriptions are available, so we use an LM judge to assess the similarity of the predicted description to the gold description, on a scale of 0 (no similarity) to 1 (very similar), in increments of 0.25. We use GPT-4.1-mini as the judge; the exact prompt can be found in Section˜A.2. Human assessment of 200 samples of the LM judge results shows 81.25% human agreement. [^8] Details of this assessment can be found in Section A.2.
Simulator score.
We also use simulator score, described in Equation˜2, which measures the Pearson correlation between the model’s true activations and its predicted activations based on the explanation $\mathcal{E}(v)$ on each sample $x$.
3.3 Baselines
To contextualize our results, we compare to two different baselines:
Nearest neighbors.
To evaluate whether explainer models exhibit nontrivial generalization to unseen features, we compare to a nearest-neighbor baseline that, given a new feature or representation, finds the single most similar feature from the training set, and outputs its associated description. Let $S_{\text{train}}$ denote the set of features used for training, and let $S_{\text{train},\ell}$ denote the subset of SAE features at layer $\ell$. We evaluate two nearest-neighbor baselines:
  1. 1.
    Layerwise nearest neighbor. We restrict retrieval to features from the same layer $\ell$, so that the retriever explanation is $\mathcal{E}_{\text{NN-layer}}(v,\ell)=D\!\left(\arg\max_{v_{i}\in S_{\text{train},\ell}}\langle v_{i},v\rangle\right)$, where $D$ is the function mapping training features to their explanations.
  2. 2.
    All-layer nearest neighbor. We retrieve from the training feature set across all layers: $\mathcal{E}_{\text{NN-all}}(v)=D\!\left(\arg\max_{v_{i}\in S_{\text{train}}}\langle v_{i},v\rangle\right)$.
SelfIE.
Following SelfIE (chen2024selfie), we directly patch $v$ into the embedding layer of a non-finetuned model, [^9] The original SelfIE paper patches representations into the third layer, but prior work has found that patching into the first few layers work about equivalently well (which we also find). Thus, we patch every activation into the embedding layer. and prompt the LM to define the feature. We use standard prompts found to work well for self-explanations (kharlapenko2024self). [^10] SelfIE prompts can be found in Section F.1. Prior works have found that SelfIE is sensitive to the scale of $v$ when it is inserted (kharlapenko2024self). We report results on the best explanation across 5 possible scales ($v$, $5v$, $10v$, $25v$, and $50v$) for each example, representing an upper bound on SelfIE performance.
We also procure gold SAE labels from Neuronpedia on the held-out SAE features. Note that the LM judge evaluates with respect to the same gold descriptions, meaning the gold SAE labels serve as a top-line number for the LM judge metric.
[…]
LM judge Simulator Simulator Simulator
Llama-3.1-8B 76.2±0.9 45.1±0.7 49.7±0.6 32.0±0.8
Llama-3.1-8B (LoRA) 76.0±0.9 45.6±0.8 50.6±0.6 34.0±0.8
Llama-3-8B 77.0±0.9 44.6±0.7 49.3±0.6 32.4±0.7
Llama-3.1-8B-Instruct 77.1±0.8 42.7±0.7 46.9±0.7 29.9±0.7
Qwen3-8B 70.3±0.9 40.6±0.8 21.1±0.7 12.30.4
Llama-3.1-70B (LoRA)
    random projection 63.9±1.0 39.5±0.8 12.6±0.4 12.2±0.4
    pre-trained proj. 74.1±0.9 45.2±0.7 33.8±0.7 20.6±0.6
NN-all 58.5±1.0 33.7±0.8 38.9±0.7 18.5±0.6
NN-layer 53.9±0.9 29.4±0.7 40.7±0.7 28.6±0.7
SelfIE Best of 5
    Llama-3.1-8B 40.1±1.0 36.4±0.8 43.3±0.6 21.0±0.6
    Llama-3.1-8B-Instr. 40.4±0.9 36.4±0.8 43.5±0.7 20.1±0.6
Gold SAE labels 100±0.0 43.3±0.8 - -
[…]
3.4 Results
We report results for our trained explainers, as well as each of the baselines, in Table˜1. Based on these results, we make several observations:
Trained explainers are effective.
For every feature type and metric, training Llama-3.1-8B to explain itself outperforms all of the baselines by a significant margin. It even outperforms the gold SAE labels when scored via the simulator, due to noise in the ground-truth feature labels themselves (see error analysis in Section˜B.5).
For explainers, matching the subject matters more than parameter count.
Among all trained models, the Llama-3.1-8B and Llama-3-8B models consistently performed the best across all four evaluations, while the Llama-3.1-8B-Instruct model was consistently a few (3-5) points behind. Meanwhile, the Qwen3-8B and the randomly-initialized Llama-3.1-70B model consistently performed the worst across all four evaluations, sometimes even underperforming baselines on the out-of-distribution evaluations. In this case, with Llama-3.1-8B as the target model, a Llama-3.1-8B explainer is significantly better than Llama-3.1-70B, showing that matching the target model matters more than a 9x increase in parameter count in this case.
Similar models are better explainers.
More generally, we observe that models closest in architecture and training regimen to Llama-3.1-8B perform the best, while models that deviate in one or both perform worse: Llama-3.1-8B and Llama-3-8B share architecture and deviate slightly in pre-training, while Llama-3.1-8B-Instruct has the same architecture but has an additional instruction-tuning step. Qwen3-8B and Llama-3.1-70B both deviate completely in either training regimen or architecture, with Qwen3-8B differing significantly in training (but having a similar architecture) and Llama-3.1-70B differing in architecture (but sharing similar pre-training regimen).
Alignment between activations improves verbalization capability.
To formalize the observation above, one way of quantifying the privileged access hypothesis is to posit that activation similarity between the explainer and target model predicts explainer performance. To investigate this, we compare the explanation capabilities of Llama-3.1-70b with both pre-trained projection and randomly initialized projection as a proxy for maximally and minimally aligned activation, respectively, given the model. As reported in Table˜1, we can recover a significant fraction of performance when we start training from the pre-trained projection, performing 15% better on SAE explanations, 2.7 times better on real activations, and 1.7 times better on activation differences.
Additionally, we quantify the activation similarity of (a subset of) models reported in Table˜1 with respect to the target model, and plot activation similarity vs. explainer performance. These plots can be found in Section˜B.2. Visually, they suggest that activation alignment is correlated with privileged access, in line with our findings above.
Taken together, these findings lead us to the following conclusion:
In feature description tasks, models’ privileged access to their internals improves explanations: they are better explained by fine-tuned versions of themselves than by other models.
Error Analysis.
Finally, we performed a qualitative error analysis of 100 samples where the LM judge scored the explainer’s prediction as 0 (details in Section˜B.5). We found that most errors stem from issues in the “gold” SAE labels from Neuronpedia (consistent with our finding that Llama-3.1-8B can outperform the gold labels), LM judge scoring inconsistencies, and inherent prediction difficulty from sparsely activating features. Only 27% of errors are attributed to genuine explainer failures.
4 Privileged Access Confers Data-Efficiency
To answer where and why privileged access may be useful, we investigate whether it confers data-efficiency advantages over alternatives. We train Llama-3.1-8B, Qwen3-8B, and the nearest neighbor baseline on subsets of the training data of various sizes and investigate scaling trends. The scaling curve for the held-out SAE features (LM judge evaluation) is reported in Figure˜5, while the scaling curves for the other metrics and evaluation settings can be found in Figure˜7.
Generally speaking, self-explanation is more data-efficient than either training another explainer model or nearest neighbors. At 0.8% of all training features (training on only 1024 samples per layer), Llama is already at 71% of its end-performance. Qwen is only at 35% of its end-performance and both forms of nearest neighbors (within- and across-layer) are at 55% of their end-performance. Note that the end-performance of Llama is already higher than both Qwen and nearest neighbors, meaning this gap is further amplified in the low data regime. Llama at 0.8% of training features achieves 92.9% of the end-performance of all-layer nearest neighbors and 77.4% of fully-trained Qwen. These results indicate that:
Training a model to generate descriptions of its own features is data-efficient, particularly compared to training other models to describe the target model’s features.
This data-efficiency underscores the practical utility of self-explanation: annotating feature descriptions is expensive, and self-explanation training allows us to recover most of the performance with orders of magnitude less data. Figure˜5 shows that retrieving the top-1 SAE feature (i.e. nearest neighbors) from a feature dictionary of 32k samples per layer performs comparably to a self-explainer model trained on only 1k samples per layer—over a hundred times fewer features.
Finally, we speculate that the small amount of required data reflects what is minimally necessary to learn consistency between activation patterns and verbalizations: perhaps minimal training is necessary if models are simply reusing existing circuits that generate activation patterns for verbalization.
Figure 5: Training self-as-explainer is more data-efficient than alternatives. We plot scaling curves of each explainer: the number of training samples per layer is plotted against explanation quality, as judged by a LM judge on held-out SAE features. We find that matching the explainer to the target (Llama-3.1-8B) is generally the most data-efficient, while using another model as explainer (Qwen) or any form of nearest neighbor is much less effective in low-data regimes, taking over 1k-10k samples per layer to beat an untrained SelfIE baseline.
5 Generalization Across Tasks
To check whether self-explanation training and the privileged access hypothesis can be extended to other domains, we run preliminary experiments for two additional explanation types: predicting the outcomes of activation patching (§2.3) and input ablation (§2.4).
5.1 Activation Patching: Training LMs to Predict the Outcomes of Activation Interventions
We examine models’ ability to explain results from activation patching experiments, which perturbs various parts of model components and examines which cause a change to the model output. This method is described in detail in Section˜2.3. We train two different explainer models—Llama-3.1-8B and Qwen3-8B—to explain each other as the target models, creating four different combinations of evaluation settings.
Training Data.
We use the CounterFact dataset (meng2022locating) as a source of counterfactual input pairs $x,x^{\prime}$, which consists of a corpus of factual subject–relation–object sentences like Paris (subject) is the capital of (relation) France (object). We prompt the model with the subject and relation and ask the model to predict the object . We then identify a counterfactual sentence $x^{\prime}$ with the same relation but differing subject and object (Rome is the capital of) and measure whether the most likely LM prediction changes (e.g. to Italy). We perform sampling to ensure an even split between changed and unchanged predictions, at each token position and layer chunk. See Section˜E.2 for details.
Baselines.
To measure the effect of fine-tuning, we prompt the pre-trained Qwen3-8B and Llama-3.1-8B models (without any explanation training) to generate explanations of the form described in Section˜2.3. The full prompts for zero-shot activation patching can be found in Section˜F.2. This is the analogous setting to the SelfIE baseline for feature description.
Metrics.
Recall from Section˜2.3 that the explainer output typically includes two parts: whether the target output would changed under intervention (changes / remains unchanged), and the content of the change (to France). Based on this, we report three different types of metrics:
  1. 1.
    Has-Changed F1: Measures correctness of the first part of the explainer’s prediction, of whether the target’s output would change under intervention. We report the Macro F1 scores over the “changed” and “unchanged” classes.
  2. 2.
    Content Match: Measures correctness of the second part of the explainer’s prediction, of whether the explainer correctly predicts the content of target model’s prediction under intervention.
  3. 3.
    Exact Match: Exact match accuracy between generated explanation and ground-truth explanation. Requires the explainer to correctly predict both parts.
Target Explainer Exact Match Has-Changed F1 Content Match
Qwen3-8B Qwen3-8B 64.0±0.4 80.2±0.3 71.0±0.4
– activation 59.9±0.4 76.5±0.4 68.6±0.4
– layer 65.0±0.4 80.0±0.3 72.7±0.4
– token 63.9±0.4 79.4±0.3 71.7±0.4
Llama-3.1-8B 54.1±0.4 73.9±0.4 65.4±0.4
Qwen-3-8B (Untrained) 5.02±0.2 24.4±0.8 18.78±0.3
Llama-3.1-8B Llama-3.1-8B 48.6±0.7 77.5±0.6 54.6±0.7
– activation 45.2±0.7 73.9±0.6 50.8±0.7
– layer 49.7±0.7 76.9±0.6 55.7±0.7
– token 47.3±0.7 76.1±0.6 53.4±0.7
Qwen3-8B 41.7±0.7 75.1±0.6 47.6±0.7
Llama-3.1-8B (Untrained) 3.2±0.1 31.4±0.4 7.0±0.2
Table 2: A target model’s activation patching outcomes are best predicted by itself. We train Qwen3-8B and Llama-3.1-8B to verbalize their own activation patching outcomes, and compare against training the other model to explain their outcomes, and also against untrained versions of themselves. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each target model and metric, bold indicates no significant difference from the best entry (paired $t$-test, $p\geq 0.05$). We also conduct experiments ablating the activation, layer, and token information from the explainer’s input. Ablating layer and token each have neglible effect, potentially because they can already be inferred from the activation.
Results
for activation patching can be found in Table˜2. We see that the privileged access hypothesis is supported: Llama-3.1-8B is best at explaining Llama-3.1-8B, while Qwen3-8B is best at explaining Qwen3-8B. All trained explainers perform better than their untrained versions, which generally guess the has-changed value randomly.
Ablations.
In addition to training each explainer model on each target model, we perform a set of ablations to isolate which component(s) of the input the explainer learned to condition on. The explainer input consists of three components: the activation $v$ to insert (– activations), the layers $\ell_{1:i}$ (– layer), and the token $t$ of the target model to patch into (– token). We train the explainer model with each aspect of the input ablated (conditioning only on the other two components) to generate the ground-truth explanations. [^11] Ablated prompts can be found in Section F.3. Our ablation studies suggest found that performance is generally robust to ablating the layer and token annotations in the input prompt. We believe this is because layer and token information is generally recoverable from the activation and input context alone. As evidence for this, we train Llama-3.1-8B to decode the layer and token from which an activation comes from, and get 98.7% accuracy (see Section˜C.1).
5.2 Input Ablations: Training LMs to Describe their Decision Rules
Similar to the activation patching experiments, we train two models—Llama-3.1-8B-Instruct and Qwen3-8B—on their own behaviors and the behaviors of each other as the target models. Using the methods described in Section˜2.4, we train models to explain how withholding or including parts of target models’ input would affect the output prediction.
Training Data.
Following recent work used to study faithful chain-of-thought (chen2025reasoningmodelsdontsay), we give $\mathcal{M}$ inputs in the form of a multiple-choice question $c$ with an injected hint $\tilde{x}$, i.e. $x=[c,\tilde{x}$], and train the explainer to predict whether $\mathcal{M}$ ’s most likely next token will change to $\tilde{x}$. Previously, chen2025reasoningmodelsdontsay found that models fail to verbalize $\tilde{x}$ in their chain-of-thought, despite $\tilde{x}$ being critical to its prediction ($\mathcal{M}([c,\tilde{x}])\neq\mathcal{M}(c)$). Thus, we investigate whether models can be trained to detect when and how their prediction will change depending on $\tilde{x}$’s presence.
We use MMLU (hendryckstest2021) as the source of multiple-choice reasoning question, and inject hints of the form $\tilde{x}$=Hint: A” to the end of the question. We randomly select $\tilde{x}$ as one of the four multiple-choice options. To make the has-changed prediction nontrivial, we construct the hint prompt such that we have a roughly equal number of samples that change and do not change according to the hint. See Section˜E.3 for details.
Metrics.
Recall from Section˜2.4 that the explainer’s prediction involves two parts, similar to activation patching: whether the target’s prediction will change when the hint is removed (e.g. output changes / output remains unchanged), and the new prediction / content of the change (e.g. to (A)). Thus, we use the same three metrics as in activation patching (§5.1).
Baselines.
Similar to Section˜5.1, we include untrained explainer baselines. The prompt for then untrained models can be found in Section˜F.3.
Target Explainer Exact Match Has-Changed F1 Content Match
Llama-Inst. Llama-Inst. 18.2±1.0 58.2±1.3 24.9±1.2
Qwen3-8B 14.6±0.9 52.4±1.2 22.9±1.1
Llama-Inst. Llama-Inst.
21.0±1.1 45.5±1.3 36.7±1.3
Qwen3-8B Qwen3-8B 62.7±1.3 73.4±1.1 70.5±1.2
Llama-Inst. 33.1±1.3 59.4±1.3 48.3±1.3
Qwen3-8B Qwen3-8B
30.2±1.2 53.2±0.5 39.9±1.3
Table 3: An explainer model can best explain its own input ablation outcomes. We train Qwen3-8B and Llama-3.1-8B to verbalize their own input ablation outcomes, and compare against training to verbalize each other’s outcomes. We also report untrained models explaining themselves. Note this table is sorted by explainer rather than target unlike previous tables. Scores (mean $\pm$ standard error) are reported for three metrics—exact match, has-changed F1, and content prediction. For each explainer model and metric, we bold all entries not significantly different from the best entry (paired $t$-test, $p\geq 0.05$). We find evidence of a different version of privileged access: models are better at explaining themselves than other models.
Results
are shown in Table˜3. We find a different pattern of privileged access in this case, as Llama-3.1-8B appears to be an overall a much weaker explainer of both itself and of Qwen. However, when the explainer (instead of target) is held constant, both models are better at explaining themselves than explaining others: Llama is better at explaining Llama than Qwen (despite being a worse explainer overall), and vice versa. One way of viewing this is that privileged access is a significant first-order term in explainer performance, but can in some cases be overcome by other factors, such as a significant gap in the task-relevant capabilities of the explainer model.
One indication of Llama’s lack of capability comes from comparison to its untrained version: while Llama can learn to better predict its own has-changed value, its content match remains around random (25%) even after training. [^12] Although untrained Llama’s content match is above random, a closer investigation reveals that it’s simply predicting the same content as the hint 81.9% of the time. This results in an above-random content match because if the model prediction remains unchanged, it typically is because the model’s original prediction matches the hint; we do not take the metric as evidence that untrained Llama is doing something non-trivial. If we only look at the has-changed F1, Llama’s ability to explain itself is on par with Qwen’s ability to explain it, despite Llama being a weaker overall explainer.
Finally, we examine untrained Qwen’s predictions and find that it a has-changed value of True only 0.4% of the time, and most of its content match comes from correctly predicting the unchanged answers (91.0% content match) compared to its changed answers (14.3% content match). This is consistent with findings on reasoning models in prior work: chen2025reasoningmodelsdontsay finds that no pre-trained LM would ever verbalize that it utilizes the hint, even when it did, similar to our finding that untrained Qwen never answers that its response changes based on the hint. This indicates that explicit fine-tuning is essential to elicit models to verbalize their true decision process.
Taken together, results on these two tasks show that:
Privileged access extends across tasks, but requires a certain model capacity; some models may be (model-agnostic) poor explainers of certain types of computation.
6 Related Work
6.1 Chain-of-Thought Faithfulness
Language models can be asked to verbalize their thought processes through chain-of-thought prompting, which offers a way for external monitoring (korbak2025chainthoughtmonitorabilitynew; baker2025monitoringreasoningmodelsmisbehavior), but prior work has found that these verbalizations can be unfaithful to their true decision-making processes (turpin2023language; lanham2023measuringfaithfulnesschainofthoughtreasoning; barez2025cot; chen2025reasoningmodelsdontsay). This line of work directly inspires our hint ablation task setup, while our work attempts to remedy the unfaithfulness through a training objective that enforces consistency between model verbalization and their actual mechanisms.
6.2 Mechanistic Interpretability Methods
The field of mechanistic interpretability sets out to (1) describe individual features and (2) connect them to construct a causal circuit that performs a certain task. While many researchers have found success doing so manually in a fine-grained way both on the neuron level (gurnee2024universal) and the circuit level for toy tasks (wang2023interpretability; nanda2023progressmeasuresgrokkingmechanistic), these methods have been challenging to scale and generalize to new models and tasks (sharkey2025open). To address scalability, automated feature description pipelines (hernandez2022natural; bills2023language; choi2024automatic; paulo2025automatically) and circuit discovery techniques (conmy2023automated; syed-etal-2024-attribution; hanna2024have; hsu2025efficient) have been introduced that require substantial computational resources. Designing a scalable interpretability method that balances correctness/robustness and efficiency is an active area of research, one of which is attribution patching that use gradients to approximate important components to the output (nanda2023attribution; syed-etal-2024-attribution).
Closely related to the current study, one family of scalable interpretability methods attempts to leverage models’ own verbalization ability without any additional training. LogitLens (nostalgebraist2020interpreting), PatchScopes (ghandeharioun2024patchscopes), SelfIE (chen2024selfie) enable models to self-interpret zero-shot, but such methods require significant hyperparameter tuning (kharlapenko2024self).  pan2024latentqa introduces Latent Interpretation Tuning, which fine-tunes models to answer questions about other models’ inputs via their hidden representations. We expand on this line of work by training models to self-verbalize their internals based on data collected from existing interpretability techniques that provide finer-grained pictures of models’ internal computations.
Finally, explanations based on special interpretability techniques are difficult for non-expert users to access, potentially requiring developers integrate many specialized interfaces (viégas2023modelusermodelexploring; chen2024designingdashboardtransparencycontrol). Our work offers an alternative path toward user understanding, in which explanations can be provided “in-band” rather than via an external interface.
6.3 Introspection & Metacognition
Recent work investigates whether models possess metacognition or introspective abilities (binder2025looking; treutlein2024connecting; laine2024memyselfaisituational; comsa2025doesmakesensespeak; plunkett2025selfinterpretabilityllmscomplexinternal). A central debate in this literature concerns whether models have privileged access to their internals, or whether introspection merely reflects their strong predictive capacity to learn external correlations (song2025language; song2025privilegedselfaccessmattersintrospection; li2025do). We add to the literature by assessing models’ introspective abilities on their own mechanisms derived from mechanistic interpretability methods.
7 Discussion: Self-Consistency as a Framework for Interpretability and Faithfulness
This paper explores improving model faithfulness and interpretability by training models to generate explanations consistent with outcomes of interpretability procedures. Across two question types, we find evidence that generated explanations take advantage of models’ privileged access to their own internal mechanisms, and on the third, we find evidence of explainer models having privileged target access to their own internals. Quantifying how factors such as explainer model capacity, target model complexity, task difficulty, etc. influence the emergence or manifestation of privileged access remains an important question for future work.
In a general sense, our results suggest that objectives that enforce self-consistency between models’ behavior and explanations may provide an avenue towards more scalable interpretability methods, faithful explanations, and alignment guarantees. The experiments in this paper derived (self-)explanation objectives from feature description, activation patching and input ablation techniques. Generalized versions of these objectives could in principle be applied to any method for generating ground-truth descriptions of model computation or behavior, e.g. to describe inferred user traits, surface adversarial triggers, or summarize influential training examples (chen2024designingdashboardtransparencycontrol; Wallace2019Triggers; pmlr-v70-koh17a).
One natural extension would also optimize for consistency in a stronger sense: rather than having fine-tuned explainer models explain static target models, enforce that the target model is always identical to the explainer model. Perhaps more interestingly, these objectives could be applied in the reverse direction, fine-tuning models so that their behavior matches generated explanations, which might be normatively correct even if inaccurate as descriptions of models themselves. Applied in this way, consistency fine-tuning could provide a new mechanism for “unsupervised alignment” without relying on large preference datasets or reinforcement learning procedures.
Acknowledgments
We would like to thank Dami Choi for technical support at early stages of this project. This work was supported by the Open Philanthropy Foundation and the MIT Quest for Intelligence. BZL is additionally funded by a Clare Boothe Luce Fellowship, and JA is funded by a Sloan Fellowship.
Appendix A Feature Descriptions: Additional Experimental Details
A.1 Explainer models
We train the following five explainer models to describe features from Llama-3.1-8B:
  1. 1.
    Llama-3.1-8B: a 32-layer, 4096-hidden-dimension, 32-attention-head model pre-trained on over 15T tokens.
  2. 2.
    Llama-3-8B: exact same network architecture as Llama-3.1-8B, with very similar training regimen (also on 15T tokens). However, Llama-3.1-8B was trained with significantly longer context length via continual pre-training, on multilingual inputs beyond just English.
  3. 3.
    Llama-3.1-8B-Instruct: took Llama-3.1-8B and did additional instruction-tuning on top of it so that it accepts chat format. We query this explainer model in a chat format.
  4. 4.
    Llama-3.1-70B: same training regimen as Llama-3.1-8B, and basically the same architecture except much larger, with 80 layers, 8192 hidden dimension, and 64 attention heads.
  5. 5.
    Qwen3-8B: it is similar to Llama-3.1-8b that they both are 4096-hidden-dimension, 32-attention-head model. but Qwen3-8b has 36 layers and a smaller MLP intermediate size (12,288 vs 14,336). Qwen is trained on approximately 36T tokens across 119 languages, while Llama is trained on 15T.
We select these models such that each model minimally (or as minimally as possible) differs from the target model in one of architecture, size, and training regimen, to investigate the effect of each of these factors.
A.2 Metric Details: LM Judge
LM Judge Prompt.
We use the following prompt for the LM judge to assess similarity between ground-truth feature descriptions and predicted feature descriptions:
  • Does this feature description accurately describe when this feature activates?
    Rate on a scale of:
    - 1 = completely unrelated to expected
    - 2 = mostly unrelated
    - 3 = somewhat related
    - 4 = related and fairly similar
    - 5 = same as expected, or highly similar (treat this as a correct match)
    If unsure between 4 and 5, choose 5.
    Examples:
    Predicted: mentions of cooking recipes
    Expected: references to financial transactions
    Correct rating: 1
    Predicted: mentions of dogs and cats
    Expected: references to farm animals
    Correct rating: 2
    Predicted: mentions of sunny weather and rain
    Expected: references to climate conditions
    Correct rating: 3
    Predicted: mentions of jazz musicians and concerts
    Expected: references to music
    Correct rating: 4
    Predicted: mentions of Shakespeares plays
    Expected: references to works by Shakespeare
    Correct rating: 5
    Now rate the following pair:
    Predicted: predicted_label
    Expected: expected_label
    Return a number from 1 to 5 and nothing else.
where predicted_label and expected_label are placeholders for the ground-truth and predicted feature descriptions.
Human Validation.
To validate our LM judge scoring methodology, we conduct a qualitative evaluation against human judgments. We sample 200 queries from the test set, each containing an (expected label, predicted label) pair, and construct 100 comparison pairs of the form $(\text{expected}_{1},\text{prediction}_{1},\text{expected}_{2},\text{prediction}_{2})$. A human annotator who has not previously seen these examples provides preference judgments on which label is superior. The annotator assigns one of three labels: 1 (prediction1 is better), 2 (prediction2 is better), or 0 (both predictions are equally good). We employ preference-based evaluation rather than absolute scoring because human annotators and the LM judge may operate with different baselines.
Our analysis reveals strong alignment between human and LM judge preferences. Out of 100 pairs, we achieve 64% exact agreement, where both the human and LM judge select the same preference. When accounting for near-agreement—weighting differences of 0.25 points as 0.75 correct and 0.5 points as 0.5 correct (e.g., cases are common where human judges rate both labels as equally good while the model assigns scores of 1.0 and 0.75)—agreement increases to 81.25%. Notably, only 3 out of 100 pairs (3%) exhibit complete divergence, where human and LM judge preferences are directly opposed (human prefers $\text{prediction}_{1}$ over $\text{prediction}_{2}$ while LM judge prefers $\text{prediction}_{2}$ over $\text{prediction}_{1}$).
Appendix B Feature Descriptions: Additional Results
B.1 Gemma-2-9B as Target Model
We replicate a subset of the evaluations in Table˜1 using Gemma-2-9B as the target model. We select Gemma-2-9B because its SAE features are also publicly available via GemmaScope (lieberum-etal-2024-gemma), and downloadable via Neuronpedia (neuronpedia). Results can be found in Table˜4.
Model SAE (LM Judge)
Gemma-2-9B 43.45%±0.91%
Gemma-2-9B-Instruct 57.12%±0.87%
Llama-3.1-8B 34.45%±0.88%
NN-all 32.40%±0.80%
NN-layer 33.07%±0.78%
SelfIE Best of 5 (Gemma-2-9B) 24.07%±0.85%
Table 4: Feature descriptions results using Gemma-2-9B as a target model. We find similar results to using Llama-3.1-8B as the target model, where self- and self-adjacent models (Gemma-2-9B, Gemma-2-9B-Instruct) are significantly better at explaining Gemma-2-9B compared to other models, top-1 nearest neighbors, and untrained SelfIE.
We find that similar results hold with Gemma-2-9B as target model as with Llama-3.1-8B as target model. Explainer models that are similar to Gemma-2-9B (Gemma-2-9B, Gemma-2-9B-Instruct) are the best at explaining Gemma-2-9B, while Llama-3.1-8B is significantly worse. Furthermore, self-explanation training outperforms nearest neighbors and untrained SelfIE, consistent to our findings in the main paper.
In the case of Gemma-2-9B, we find that instruction-tuning makes a significant difference for explanation ability. Even though the Gemma-2-9B-Instruct model may be slightly unaligned from the Gemma-2-9B base model thanks to instruction-tuning, this gap matters less in this case than additional abilities conferred to the model via instruction tuning.
B.2 Quantifying privileged access via activation alignment
Let $h^{\mathcal{E}}_{\ell,t}(x)$ denote activations from the explainer model $\mathcal{E}$ at layer $\ell$, token position $t$, and $h^{\mathcal{M}}_{\ell,t}(x)$ denote activations from the target model $\mathcal{M}$ at layer $\ell$, token position $t$. We consider two ways of measuring activation similarity:
  1. 1.
    Dot-product similarity: We measure the dot-product similarity between activations from the explainer vs. target models on the same inputs at the same locations:
    $s(\mathcal{E},\mathcal{M})=\mathbb{E}_{x,\ell,t}\big[\left\langle h^{\mathcal{E}}_{\ell,t}(x),h^{\mathcal{M}}_{\ell,t}(x)\right\rangle\big]$
  2. 2.
    SAE activation pattern similarity: Using a similar approach to the simulator correlation equation 2, for each feature $v$, we examine the activation pattern between each model on the top exemplars $x$ where $v$ activated the most. [^13] Top exemplars are downloaded from Neuronpedia. We then measure the Pearson correlation coefficient between the models’ activation patterns.
    $s(\mathcal{E},\mathcal{M})=\mathbb{E}_{\ell,v}\mathbb{E}_{x\mid v}\left[\texttt{corr}_{t}\left(\left\langle h^{\mathcal{E}}_{\ell,t}(x),v\right\rangle,\left\langle h^{\mathcal{M}}_{\ell,t}(x),v\right\rangle\right)\right]$
For each similarity metric and each explainer model (fixing 3.1-8B as the target model), we provide raw similarity values in Table˜5. We then plot explainer similarities vs. explainer performance for each of the four (metric, feature type) pairings, as described in Section˜3.2. Plots can be found in Figure˜6. We find that explainer performance tracks the activation similarity of the explainer, indicating alignment between activations predicts verbalization capability.
Model SAE Act. Pattern Similarity Dot Product Similarity
Llama-3-8B $93.66\%\pm 14.89\%$ $96.61\%\pm 1.73\%$
Llama-3.1-8B-Instruct $88.63\%\pm 18.3\%$ $85.27\%\pm 5.9\%$
Llama-3.1-70B (pre-trained projection) $54.25\%\pm 14.57\%$ $21.48\%\pm 4.53\%$
Qwen3-8B $0.10\%\pm 11.94\%$ $0.10\%\pm 0.49\%$
Table 5: Similarity between each explainer model and Llama-3.1-8B activations, measured by activation pattern similarities between SAE features and dot-product similarities between activations. The SAE-token-level correlations between each explainer model and Llama-3.1-8B is measured with same correlation metric as the simulator correlation. Dot-product similarity is taken directly between the two models’ activations at layer $\ell$, token position $t$. We report averages and standard deviations.
[…]
This correlation holds on all metrics: activation alignment is generally positively correlated with explainer quality, with the exception of a few outliers (e.g. Llama-3.1-70B is good at explaining held-out SAE features based on simulator scores). This solidifies our findings in Section˜3.4: alignment between activations contributes to verbalization capability.
B.3 Additional Scaling Curves
We plot scaling curves for Llama-3.1-8B, Qwen3-8B, and nearest neighbors as explainer models for the Llama-3.1-8B target model. After training on various amounts of training data, we plot the simulator scores for each of the three feature types described in Section˜3.2. These plots can be found in Figure˜7.
[…]
We find that the general trend of self-explanation (Llama-3.1-8B) being significantly more data efficient than explaining with another model (Qwen-3-8B) or using nearest neighbors holds on all three types of features when scored using simulator correlations, similar to our findings using LM judge on SAE features in Section˜4.
B.4 Is Layer Annotation Necessary to Describe SAE Features?
We investigate whether the explainer model relies on layer-specific information when generating feature descriptions. Specifically, we examine whether features extracted at layer $\ell$ require knowledge of that layer annotation or perform specialized computations at that particular layer.
Our experiments reveal that layer information plays a surprisingly minimal role in the model’s description process. When we remove the layer from the input “What does feature $v$ mean at layer $\ell$?” or replace it with an incorrect layer number, the generated descriptions remain largely unchanged. Across 100 test set prompts with a fixed rollout of 10 tokens, the ablated generations achieve 77.3% and 75.2% token-level matches with the original outputs, respectively. These surprisingly high match rates are conservative estimates, as they do not account for semantically equivalent responses that differ only in token ordering or minor phrasing variations.
These findings potentially suggest that the model does not need to distinguish which layer an SAE feature originates from to complete the task successfully, provided that SAEs across layers map onto roughly similar latent spaces. This interpretation is further corroborated by the results in Table˜1, where nearest neighbor retrieval across all SAE layers outperforms retrieval restricted to the same layer for the feature description task.
B.5 Error Analysis
We perform a human qualitative analysis on 100 test set outputs from the explainer model that received a score of 0 from the LM judge in Table˜1. Our analysis reveals that sources of error extend beyond the explainer model itself and can be attributed to multiple stages of the evaluation pipeline, including (1) the original feature labels generated by the LM based on activation exemplars and (2) the validity of the LM judge’s assessment when comparing the explainers’ output against the original labels. Errors from these sources do not reflect failures of the explainer model.
We categorize the 100 low-scoring cases into five distinct categories, as illustrated in Figure 8. For original feature label errors, we distinguish between noisy/ambiguous labels and incorrect labels, which respectively partially and fully miscategorize the activation exemplars; in some cases, the explainer models provide better labels than the original auto-labeling pipeline. For example, consider a feature that activates on patterns like “rather than…” and “not…”. [^14] https://www.neuronpedia.org/llama3.1-8b/9-llamascope-res-131k/96078 The Neuronpedia “gold” label incorrectly describes it as “technical terms and phrases related to computational or algorithmic processes,” whereas our fine-tuned explainer correctly identifies it as “phrases that indicate comparisons or contrasts in ideas or concepts.”
LM judge error cases occur when the LM judge mistakenly considers the LM output as incorrect due to slight differences between the expected and predicted labels. The final two categories are both explainer model errors (Explainer model error and Sparse activation), but Sparse activation features have extremely rare activation patterns ($0.001\%$) that are inherently difficult to characterize due to insufficient examples.
Critically, only 27.7% of cases scored 0 by the LM judge represent genuine, unexplainable errors from the finetuned explainer, which suggests that automated metrics may underestimate the explainer model’s true performance.
Figure 8: Pie chart of human qualitative labelling of 100 exemplars with a 0 score from the LM judge. It reveals that only 27% are genuine explainer model errors, with the remaining errors attributable to low-quality autointerp labels (43.5%), LM judge error (11%), and inherent prediction difficulty due to sparse activation (18%).
Appendix C Activation Patching: Additional Results
C.1 Ablation Analysis
We analyze why ablating the layer and token in Table˜2 makes very little difference in the final result. We train Llama-3.1-8B to decode the layer and token information from the input context and activation value alone, i.e. given the prompt
[…]
we train the model to respond with form
[…]
We use the same train/test set division as for the main set of experiments in Section˜5.1, and get 98.7% exact match accuracy on the test set.
Appendix D Explainer Model Prompts
The list of explainer model prompts and output formats for each explanation type can be found below.
Feature Descriptions.
[…]
Activation Patching.
Given any of the following prompts:
[…]
The explainer model is trained to generate one of the following explanations, depending on whether the patched output is different from the original output:
[…]
Input Ablation.
Given the following prompt,
  • [SYSTEM]
    The following are multiple choice questions (with a correct answer). Output only the answer letter (A, B, C, or D) and nothing else, in the format Answer: x, where x is one of A, B, C, or D.
    [USER]
[…]
    If the hint were removed how would the assistant answer change?
    [ASSISTANT]
where [SYSTEM], [USER], and [ASSISTANT] are role tokens for the chat template (we exclusively use chat models to evaluate this setting). The model is trained to predict one of the following outputs, depending on if the hint changes the model prediction, $\mathcal{M}(c,\tilde{x})\neq\mathcal{M}(c)$:
[…]
Appendix E Data Generation Pipeline Details
Below, we include additional details about how we created our training (and test) datasets, as well as dataset statistics for each dataset.
E.1 Feature Descriptions
We download SAE features from LlamaScope (he2024llamascope), with explanations from Neuronpedia (neuronpedia). We use features from Llama-3.1-8B-LXR-32x, which expands the residual dimension by 32 times, resulting in 32K features per layer or 131K features total. We hold out 50 features per layer for testing. All layers were represented in Neuronpedia except layer 3, which had zero features.
E.2 Activation Patching
The activation patching data was created from counterfactual pairs from CounterFact (meng2022locating). The main data creation procedure is described in Section˜2.3. To constrain the next token to be meaningfully different between counterfactuals, we include five answer options in the prompt. The answer options are the same between counterfactual prompts. An example of a pair of counterfactual inputs can be found below:
  • Rome is the capital of
    Respond with one of France or Spain or UK or Italy or Egypt or unknown and nothing else.
    Paris is the capital of
    Respond with one of France or Spain or UK or Italy or Egypt or unknown and nothing else.
To create the activation patching dataset, we perform a comprehensive search over all layer chunks (in increments of 8 for Llama and 9 for Qwen) and token positions. To reduce the risk of the model picking up on spurious correlations, we filter the dataset to ensure roughly equal representation across each has-changed category while avoiding over-representation of any specific (token, layer) combination. Without this balancing, a majority of the has-changed labels would be False, except on certain token-layer combinations where the has-changed label would always be True—it would be incredibly easy for the model to perform well by simply picking up on these surface-level correlations and predicting the majority label.
A statistics report of the data showing the number of samples for each layer, token type, and has-changed label can be found in Table˜6 for the Llama-3.1-8B target model as an example. We divide the prompt into several token types, which are as follows:
  1. 1.
    Subject Final refers to the final token of subject, e.g. Rome in the above example.
  2. 2.
    Relation refers to any of the relation tokens, e.g. any of is the capital of.
  3. 3.
    Answer Options refers to any of the multiple-choice answers in the prompt. The Orig Answer Option is the answer option corresponding to $x$, e.g. Italy, while the New Answer Option is the answer option corresponding to $x^{\prime}$, e.g. France. The Other Answer Option is all of the other answer options, e.g. Spain, UK, Egypt, and unknown.
  4. 4.
    For all of the other tokens in the dataset, we take the token itself as its own category.
Data Split Token Type Layer Has-Changed
False True
Train Changed Answer Option (0, 1, 2, 3, 4, 5, 6, 7) 0 104
(8, 9, 10, 11, 12, 13, 14, 15) 0 35
(16, 17, 18, 19, 20, 21, 22, 23) 0 600
(24, 25, 26, 27, 28, 29, 30, 31) 0 600
Orig Answer Option (0, 1, 2, 3, 4, 5, 6, 7) 132 308
(8, 9, 10, 11, 12, 13, 14, 15) 52 1266
(16, 17, 18, 19, 20, 21, 22, 23) 126 128
(24, 25, 26, 27, 28, 29, 30, 31) 133 16
Other Answer Option (0, 1, 2, 3, 4, 5, 6, 7) 634 195
(8, 9, 10, 11, 12, 13, 14, 15) 693 183
(16, 17, 18, 19, 20, 21, 22, 23) 539 269
(24, 25, 26, 27, 28, 29, 30, 31) 541 63
Relation (0, 1, 2, 3, 4, 5, 6, 7) 792 911
(8, 9, 10, 11, 12, 13, 14, 15) 806 149
(16, 17, 18, 19, 20, 21, 22, 23) 826 73
(24, 25, 26, 27, 28, 29, 30, 31) 792 55
Subject Final (0, 1, 2, 3, 4, 5, 6, 7) 102 1101
(8, 9, 10, 11, 12, 13, 14, 15) 101 448
(16, 17, 18, 19, 20, 21, 22, 23) 130 373
(24, 25, 26, 27, 28, 29, 30, 31) 114 64
Test Changed Answer Option (0, 1, 2, 3, 4, 5, 6, 7) 0 72
(8, 9, 10, 11, 12, 13, 14, 15) 0 17
(16, 17, 18, 19, 20, 21, 22, 23) 0 200
(24, 25, 26, 27, 28, 29, 30, 31) 0 200
Orig Answer Option (0, 1, 2, 3, 4, 5, 6, 7) 53 109
(8, 9, 10, 11, 12, 13, 14, 15) 28 545
(16, 17, 18, 19, 20, 21, 22, 23) 47 55
(24, 25, 26, 27, 28, 29, 30, 31) 71 5
Other Answer Option (0, 1, 2, 3, 4, 5, 6, 7) 282 100
(8, 9, 10, 11, 12, 13, 14, 15) 297 88
(16, 17, 18, 19, 20, 21, 22, 23) 226 109
(24, 25, 26, 27, 28, 29, 30, 31) 238 21
Relation (0, 1, 2, 3, 4, 5, 6, 7) 294 397
(8, 9, 10, 11, 12, 13, 14, 15) 335 54
(16, 17, 18, 19, 20, 21, 22, 23) 342 20
(24, 25, 26, 27, 28, 29, 30, 31) 348 22
Subject Final (0, 1, 2, 3, 4, 5, 6, 7) 29 449
(8, 9, 10, 11, 12, 13, 14, 15) 53 211
(16, 17, 18, 19, 20, 21, 22, 23) 58 148
(24, 25, 26, 27, 28, 29, 30, 31) 56 21
Table 6: Dataset statistics for the activation patching data on the Llama-3.1-8B target model.
E.3 Input Ablations
The input ablations data was created by injecting hints into MMLU questions (hendryckstest2021) following chen2025reasoningmodelsdontsay.
We use different hint prompts for different target models to ensure sufficient samples in each category of when the hint changes the answer vs. doesn’t change the answer. For Llama-3.1-8B-Instruct, our hint $\tilde{x}$ has format
For Qwen-3-8B, our hint $\tilde{x}$ has format
  • Hint: The answer is B
The dataset breakdown of changed/unchanged for each test split and each target model can be found in Table˜7.
Target Model Data Split Changed Unchanged
Llama-3.1-8B-Instruct Train 4739 7903
Test 518 877
Qwen3-8B Train 8517 4125
Test 929 466
Table 7: Dataset statistics for the input ablations data on each of the Llama-3.1-8B-Instruct and Qwen-3-8B target models and each split.
Appendix F Untrained Model Prompts
As baselines for all three explanation tasks, we prompt an untrained model to explain the effect of their activations. For feature descriptions, we derive prompts from prior literature on zero-shot self-explanations (kharlapenko2024self). Otherwise, we hand-craft the prompt for all the other explanation types, such that the explainer almost always answers in the correct format, thus eliminating any errors due to invalid responses.
F.1 Feature Descriptions (SelfIE)
The untrained baseline for feature descriptions is SelfIE (chen2024selfie). The following are prompts that we use for SelfIE, which we adapted from kharlapenko2024self. For base models we use
[…]
and for instruct models we use
[…]
F.2 Activation Patching
For activation patching, we use prompts of the following form:
[…]
    Respond with exactly one of the two options below, and nothing else:
    The output would remain unchanged from <<<X>>>.
    The output would change to <<<X>>>.
    Replace X with the next token of the answer following the input. Make sure X is a *single* next token.
    Example Outputs:
    The output would remain unchanged from <<<Paris>>>.
    The output would change to <<<doctor>>>.
    The output would remain unchanged from <<<Apple>>>.
    The output would change to <<<unknown>>>.
F.3 Input Ablations
For input ablations, we use prompts of form
  • [SYSTEM]
    The following are multiple choice questions (with a correct answer). Output only the answer letter (A, B, C, or D) and nothing else, in the format Answer: x, where x is one of A, B, C, or D.
    [USER]
[…]
    If the hint were removed how would the assistant answer change?
    Respond with exactly one of the two options below, and nothing else:
    The output would remain unchanged from <<<Answer: X>>>.
    The output would change to <<<Answer: X>>>.
    Replace X with the answer letter (A, B, C, or D).
    Do not include any explanations, reasoning, or extra text.
    Example outputs:
    The output would remain unchanged from <<<Answer: C>>>.
    The output would change to <<<Answer: B>>>.